Current time is about  1:05 pm,  and the large hand and the small hand on a clock have the same direction.
When such an event will happen for the next time?  In other words,  when the large  (minute)  hand and the small  (hour)  hand will have the same direction the next time?

Answer by greenestamps(13216)



The two hands point in the same direction 11 times every 12 hours, because the minute hand makes 12 revolutions in 12 hours while the hour hand makes only 1.
So the two hands point in the same direction every 12/11 hours.

The two hands point in the same direction at 12:00, so the time shortly after 2:00 when the two hands point in the same direction is 24/11 hours after 12:00.

24/11 = 2 2/11 hours
2/11 hours = 120/11 = 10 10/11 minutes
10/11 minutes = 600/11 = 54 6/11 seconds
6/11 = 0.545454...

The time will be 2:10:54.55